OpenStudy (anonymous):

solve y=x^2-4x+3 y=2x-2

OpenStudy (blank ):

There is two solutions: x=1,y=0 x=5,y=8

OpenStudy (blank ):

(1, 0) and (5, 8)

OpenStudy (blank ):

Tell me if you need the steps.

OpenStudy (anonymous):

yes please i would like to understand this more

OpenStudy (blank ):

Just a second...

OpenStudy (blank ):

y=x^(2)-4x+3 y=2x-2 Since the equation was solved for y, replace all occurrences of y in the other equations with the solution (x^(2)-4x+3). y=x^(2)-4x+3 x^(2)-4x+3=2x-2 Move all terms not containing x to the right-hand side of the equation. y=x^(2)-4x+3_ x^(2)-6x+3=-2 To set the left-hand side of the equation equal to 0, move all the expressions to the left-hand side. y=x^(2)-4x+3 x^(2)-6x+5=0 Factor the trinomial x^(2)-6x+5. y=x^(2)-4x+3 (x-1)(x-5)=0 Set each of the factors of the left-hand side of the equation equal to 0. y=x^(2)-4x+3 x-1=0 x-5=0 Since -1 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 1 to both sides. y=x^(2)-4x+3 x=1 x-5=0 Set each of the factors of the left-hand side of the equation equal to 0. y=x^(2)-4x+3 x=1 x-5=0 Since -5 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 5 to both sides. y=x^(2)-4x+3 x=1 x=5 The complete solution is the set of the individual solutions. y=x^(2)-4x+3 x=1,5 Expand the exponent (2) to the expression. For x=1 y=(1)-4(1)+3 Multiply -4 by each term inside the parentheses. For x=1 y=(1)-4+3 Add 3 to -3 to get 0. For x=1 y=0 Squaring a number is the same as multiplying the number by itself (5*5). In this case, 5 squared is 25. For x=5 y=(25)-4(5)+3 Multiply -4 by each term inside the parentheses. For x=5 y=(25)-20+3 Add 3 to 5 to get 8. For x=5 y=8 The final solution is the set of all solutions. x=1,y=0 x=5,y=8 Hope this helps. I used the Substitution method here.

OpenStudy (anonymous):

thank you this helped alot